OpenStudy (anonymous):

Verify the following identity: (1-cos2θ)/(sin2θ) = tanθ

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

sin2x = 2sinxcosx cos2x = cos^2 - sin^2 1-cos2x = sin^2 +(1-cos^2) = 2sin^2 sin2x = 2sinxcosx \[\frac{2\sin ^{2}x}{2\sin x \cos x} = \frac{\sin x}{\cos x}=\tan x\]